OpenStudy (anonymous):

Test to see if series is convergent or divergent? If convergent find the sum??

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ (1 + 2^{n} )}{ (3^{n}) }\]

OpenStudy (anonymous):

any help is appreciated...

OpenStudy (sirm3d):

\[\large \sum_{n=0}^{+\infty}\left( \frac{ 1 }{ 3^n }+\frac{ 2^n }{ 3^n } \right)\] \[\large =\sum_{n=0}^{+\infty} \left( \frac{ 1 }{ 3 } \right)^n+\sum_{n=0}^{+\infty} \left( \frac{ 2 }{ 3 } \right)^n\] BOTH look like infinite geometric series to me.